3.a. Define elasticity of supply and find the price from the given statement: If Es of a good is 2 and a firm supplies 200 units at price of Rs 8 per unit, then at what price will the firm supply 250 units.


3.b. Calculate the elasticity of supply if a 15 %increase in the price of soya bean oil increases its supply from 300 to 345 units.


1
Expert's answer
2021-04-06T18:42:10-0400

3.a. The price elasticity of supply can be found as follows:


"E_s=\dfrac{\%\Delta Q}{\%\Delta P}=\dfrac{\dfrac{Q_2-Q_1}{0.5(Q_2+Q_1)}}{\dfrac{P_2-P_1}{0.5(P_2+P_1)}},""\dfrac{P_2-P_1}{0.5(P_2+P_1)}=\dfrac{\dfrac{Q_2-Q_1}{0.5(Q_2+Q_1)}}{E_s},""\dfrac{P_2-8}{0.5(P_2+8)}=\dfrac{\dfrac{250-200}{0.5(250+200)}}{2},""\dfrac{P_2-8}{0.5(P_2+8)}=0.11,""P_2=Rs\ 8.93."

3.b.


"E_s=\dfrac{\%\Delta Q}{\%\Delta P}=\dfrac{\dfrac{Q_2-Q_1}{0.5(Q_2+Q_1)}}{\%\Delta P},""E_s=\dfrac{\%\Delta Q}{\%\Delta P}=\dfrac{\dfrac{345-300}{0.5(345+300)}}{0.15}=0.93"